OpenStudy (aivantettet26):

If two thermometers, one reading Celsius and the other Kelvin, are inserted in the same system, under what circumstance will they both have the same reading? What will be the system's temperature when the absolute thermometer reads twice the numerical reading of the Celsius thermometer?

OpenStudy (mrnood):

There is always a fixed offset of approx. 273 K between temperature in Celsius and temperature in K So can they ever be the same?

OpenStudy (mrnood):

For the second part you can see that K= C+273 so if the value in K = 2C then 2C= C+273 Solve for C

OpenStudy (aivantettet26):

@MrNood that's what I thought. C and F has -40. but C and K? Although I read something about 0 K and -273.18 C to be the absolute freezing point of all molecules. But I'm not sure if that's the answer the book is looking for.

OpenStudy (aivantettet26):

@MrNood I'm getting 273.18

OpenStudy (mrnood):

0K = -273.18 C These are not numerically equal!

OpenStudy (aivantettet26):

@MrNood I'm getting 273.18